OpenStudy (freemap):

Graph of y + 1 = 1/3 (x –6)?

OpenStudy (freemap):

please help

OpenStudy (joannablackwelder):

Can you solve for y?

OpenStudy (freemap):

like 1/3x + 2 y=1/3+3 I don't know

OpenStudy (joannablackwelder):

That is a good try, but when we distribute the 1/3, I get 1/3x-2 instead. Does that make sense?

OpenStudy (freemap):

yes, because of -6

OpenStudy (joannablackwelder):

Right, good, so that gives us y+1=1/3x-2 so far

OpenStudy (joannablackwelder):

Then to move the +1 over to the other side we need to subtract 1 on both sides: